OpenStudy (anonymous):

The current in a solenoid is changed from 1.1 A to 4.4 A. By what factor does the magnitude of the magnetic field inside the solenoid increase? A. 2 B. 4 C. 0.5 D. 0.25

OpenStudy (michele_laino):

let's suppose that our solenoid has N turns and its length is L, then the field H inside it is given by the subsequent formula: \[\Large H = {\mu _0}\frac{{NI}}{L},\quad {\mu _0} = 12.56 \times {10^{ - 6}}Henry/meter\]

OpenStudy (anonymous):

1.256E-5 ?

OpenStudy (michele_laino):

yes! so we can write: \[\Large \begin{gathered} {H_1} = {\mu _0}\frac{{N{I_1}}}{L},\quad {H_2} = {\mu _0}\frac{{N{I_2}}}{L} \hfill \\ \hfill \\ \frac{{{H_2}}}{{{H_1}}} = \frac{{{I_2}}}{{{I_1}}} \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

ok! what do we plug in?

OpenStudy (michele_laino):

and therefore: \[\Large \frac{{{H_2}}}{{{H_1}}} = \frac{{{I_2}}}{{{I_1}}} = \frac{{4.4}}{{1.1}} = ...\]

OpenStudy (anonymous):

4! so choice B (4) is our solution? :O

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

yay!! thanks!!:D

OpenStudy (michele_laino):

:)

Directrix (directrix):

@Daniellelovee